# RoboPEPP: Vision-Based Robot Pose and Joint Angle Estimation through Embedding Predictive Pre-Training

### v1

@@ Abstract
Vision-based pose estimation of articulated robots with unknown joint angles has applications in collaborative robotics and human-robot interaction tasks. Current frameworks use neural network encoders to extract image features and downstream layers to predict joint angles and robot pose. While images of robots inherently contain rich information about the robot’s physical structures, existing methods often fail to leverage it fully; therefore, limiting performance under occlusions and truncations. To address this, we introduce RoboPEPP, a method that fuses information about the robot’s physical model into the encoder using a masking-based self-supervised embedding-predictive architecture. Specifically, we mask the robot’s joints and pre-train an encoder-predictor model to infer the joints’ embeddings from surrounding unmasked regions, enhancing the encoder’s understanding of the robot’s physical model. The pre-trained encoder-predictor pair, along with joint angle and keypoint prediction networks, is then fine-tuned for pose and joint angle estimation. Random masking of input during fine-tuning and keypoint filtering during evaluation further improves robustness. Our method, evaluated on several datasets, achieves the best results in robot pose and joint angle estimation while being the least sensitive to occlusions and requiring the lowest execution time.
@@ 3.2.3 Loss Functions
where $\Phi_{pred}$ and $\Phi_{gt}$ represent the predicted and ground truth joint angles, respectively. To enhance training convergence, mean-variance normalization is applied to $\Phi_{gt}$. For keypoint detection, we utilize the focal loss [21]:
$\mathcal{L}_{kp}=\text{FocalLoss}(H_{pred},H_{gt})$ (3)
where $H_{\text{pred}}$ and $H_{\text{gt}}$ denote the predicted and ground truth heatmaps, respectively. $H_{\text{gt}}$ is generated using unnormalized Gaussian probability density functions centered at each keypoint with a 2-pixel standard deviation.
The overall training loss is a weighted combination of the two losses: $\mathcal{L}=\mathcal{L}_{joint}+\alpha(t)\mathcal{L}_{kp}.$, where $\alpha(t)$, dependent on epoch $t$, balances their relative importance. Since the joint angles are predicted in radians, $\mathcal{L}_{joint}$ tends to be much smaller than $\mathcal{L}_{kp}$, especially in early training. To address this, $\alpha(t)$ is initialized at 0.0001, increased to 0.01 after 5 epochs, 0.1 after 10 epochs, and finally to 1 after 40 epochs, ensuring a balanced curriculum for training.
@@ 3.3 Robot Pose Estimation
Keypoint Filtering: The final layer of the Keypoint Net contains a sigmoid nonlinearity, that produces heatmaps with pixel values between 0 and 1, representing keypoint confidence at each pixel. The pixel with the highest confidence indicates the keypoint location. However, when only a portion of the robot is visible, some keypoints may lie outside the image, leading to low confidence scores across the heatmap for these keypoints (Fig. 4). Selecting the pixel with the highest confidence in such cases can be misleading, as no pixel accurately represent the true keypoint. To address this, during evaluation, we apply a threshold, only considering keypoints with confidence above it. For use with a PnP algorithm [19] for pose estimation, we require a minimum of four 2D-3D correspondences. If fewer than four keypoints remain after filtering, we iteratively reduce $\epsilon$ by 0.025 until at least four keypoints are retained.
Figure 4: The examples show predicted heatmaps for Joint 5 and the End-Effector overlaid on the original image. The End-Effector, being positioned outside the field of view, produces noisy heatmaps with lower confidence (measured by peak values). Heatmap pixel values are normalized here for visual clarity.
@@
where $P_{pred}^{i}$ and $P_{proj}^{i}$ are the $i^{th}$ keypoint in $P_{pred}$ and $P_{proj}$, respectively. During sim-to-real finetuning, we use a lower learning rate than in the prior supervised training. Further, to prevent model collapse, the Keypoint Net’s learning rate is set close to zero.
@@
Region of Interest (RoI) Detection: During evaluation, we utilize the GroundingDINO [22] object detection model to automatically locate the region of interest around the robot. The detected region is cropped and resized to $224\times 224$ pixels. During training, we use ground truth bounding box information. However, to ensure our model’s robustness to region-of-interest detection, we employ a training curriculum: we use the ground truth bounding boxes and expand them by adding random offsets sampled from the uniform distribution $\mathcal{U}(0,\lambda)$ to their edges, with $\lambda$ progressing from 0 (first 30 epochs) to 30 pixels at epoch 30, 50 pixels at epoch 50, 80 pixels at epoch 70, 100 pixels at epoch 90, and 120 pixels at epoch 110. This approach encourages the model to generalize effectively, even with noisy region-of-interest detection during inference.
@@ 4.2.1 Robot Pose Prediction
RoboPEPP yields the highest scores across all sequences (except for Kuka DR where it remains competitive) among methods with unknown joint angles and bounding boxes. HPE [5], on the other hand, shows sensitivity to bounding box selection, with its performance dropping when ground truth bounding boxes are unavailable. Our analysis has shown that using bounding boxes just 5 pixels wider than ground truth reduces HPE’s accuracy by up to 25% on the Panda Photo test set and by around 50% with 10-pixel wider boxes. While both RoboPEPP and HPE are trained using ground truth bounding boxes, RoboPEPP’s training strategy (Sec. 3.3) reduces its dependency on them. Notably, RoboPEPP outperforms HPE on most sequences even when HPE has acces to known bounding boxes during evaluation.
@@ 4.2.4 Percentage of Correct Keypoints
The accuracy of 2D keypoint detection affects the overall performance of RoboPEPP. Therefore, in Table 4, we report the percentage of correct keypoints (PCK) within thresholds of 2.5, 5, and 10 pixels on the Panda Photo dataset. Since HPE [5] and RoboPose [16] do not rely on 2D keypoint detection for pose estimation, we include only DREAM [18] as a baseline for comparison. RoboPEPP achieves high average PCK scores, with 0.43 @ 2.5 pixels, 0.73 @ 5 pixels, and 0.95 @ 10 pixels. While DREAM outperforms RoboPEPP on certain metrics, such as PCK@2.5 pixels on the Panda AK and Panda XK, RoboPEPP demonstrates superior accuracy across most other metrics and on average, highlighting its robust keypoint detection performance.
@@ 5 Conclusion
Sim-to-Real Self-Supervised Training: To bridge the sim-to-real gap, the trained networks are fine-tuned on real datasets with self-supervised training, as described in Sec. 3.3. An AdamW optimizer [23] is used with learning rates of $10^{-7}$ for the encoder and predictor, $10^{-5}$ for the joint network. The learning rate for the keypoint network is set close to zero to prevent model collapse. These learning rates are decreased by a factor of $10^{8}$ over the training process. Models are fine-tuned separately for each real-world dataset for 10 epochs with a batch size of 64.
@@
We utilize the pre-trained GroundingDINO [22] object detection model to identify the region of interest, as described in Sec. 3.3. GroundingDINO is a highly accurate open-set object detector that accepts an (image, text) pair as input and outputs bounding boxes corresponding to regions of the image described by the text query. For all real and photo-realistic test datasets, we use the text query “robotic arm.” However, for the Panda, Kuka, and Baxter domain-randomized datasets, we use the query “robot” because these images often contain multiple objects, some of which resemble arms and can confuse the detection model. All other parameters of GroundingDINO are left at their default values. To address scenarios where only a portion of the robot is detected, we expand all the detected bounding boxes, especially for real datasets. Increasing all the bounding box sizes by 100 pixels on all sides generally yields robust robot pose estimation results. However, some fine-tuning of this parameter may be necessary for optimal performance depending on the specific dataset. Nonetheless, high performance is obtained even without fine-tuning.
@@
We evaluate our method on the DREAM dataset [18], which includes sequences from three robots: Franka Emika Panda (Panda), Kuka iiwa7 (Kuka), and Rethink Robotics Baxter (Baxter). The dataset provides training and testing sequences in both synthetic and real-world settings, as detailed in Table A2. The synthetic data comprises domain-randomized (DR) and photo-realistic (Photo) sequences. For real-world data, sequences of the Panda robot were captured using Microsoft Azure Kinect (AK), Xbox 360 Kinect (XK), and Intel RealSense D415 (RS) cameras, with the cameras positioned at fixed locations. Additionally, the Panda ORB dataset was collected using a RealSense camera but with varying camera placements. Example images from each dataset sequence are illustrated in Fig. A1.

### v2

@@ Abstract
Vision-based pose estimation of articulated robots with unknown joint angles has applications in collaborative robotics and human-robot interaction tasks. Current frameworks use neural network encoders to extract image features and downstream layers to predict joint angles and robot pose. While images of robots inherently contain rich information about the robot’s physical structures, existing methods often fail to leverage it fully; therefore, limiting performance under occlusions and truncations. To address this, we introduce RoboPEPP, a method that fuses information about the robot’s physical model into the encoder using a masking-based self-supervised embedding-predictive architecture. Specifically, we mask the robot’s joints and pre-train an encoder-predictor model to infer the joints’ embeddings from surrounding unmasked regions, enhancing the encoder’s understanding of the robot’s physical model. The pre-trained encoder-predictor pair, along with joint angle and keypoint prediction networks, is then fine-tuned for pose and joint angle estimation. Random masking of input during fine-tuning and keypoint filtering during evaluation further improves robustness. Our method, evaluated on several datasets, achieves the best results in robot pose and joint angle estimation while being the least sensitive to occlusions and requiring the lowest execution time. The code is available at https://github.com/raktimgg/RoboPEPP.
@@ 3.2.3 Loss Functions
where $\Phi_{pred}$ and $\Phi_{gt}$ represent the predicted and ground truth joint angles, respectively. To enhance training convergence, mean-variance normalization is applied to $\Phi_{gt}$. For keypoint detection, we utilize the focal loss [22, 18] ($\mathcal{L}_{kp}$):
$\frac{-1}{\mathcal{S}}\sum_{i=1}^{\mathcal{H}}\sum_{j=1}^{\mathcal{W}}\begin{cases}(1-p_{ij})^{\alpha}\log(p_{ij}),&\text{if }y_{ij}=1,\\ (1-y_{ij})^{\beta}(p_{ij})^{\alpha}\log(1-p_{ij}),&\text{otherwise}\end{cases}$ (3)
with $p_{ij}$ and $y_{ij}$ being the pixel values at $(i,j)$ of the predicted $(H_{\text{pred}})$ and ground truth $(H_{\text{gt}})$ heatmaps, respectively; $\{\alpha,\beta\}=\{2,4\}$ balance the loss terms; $\mathcal{S}$ is number of pixels with $y_{ij}=1$ ($\mathcal{S}=1$ if none exist). $H_{\text{gt}}$ is created using unnormalized Gaussian probability density functions centered at each keypoint with a 2-pixel standard deviation.
The overall training loss is a weighted combination of the two losses: $\mathcal{L}=\mathcal{L}_{joint}+\alpha(t)\mathcal{L}_{kp}.$ where $\alpha(t)$, dependent on epoch $t$, balances their relative importance. Since the joint angles are predicted in radians, $\mathcal{L}_{joint}$ tends to be much smaller than $\mathcal{L}_{kp}$, especially in early training. To address this, $\alpha(t)$ is initialized at 0.0001, increased to 0.01 after 5 epochs, 0.1 after 10 epochs, and finally to 1 after 40 epochs, ensuring a balanced curriculum for training.
@@ 3.3 Robot Pose Estimation
Keypoint Filtering: The final layer of the Keypoint Net contains a sigmoid nonlinearity, that produces heatmaps with pixel values between 0 and 1, representing keypoint confidence at each pixel. The pixel with the highest confidence indicates the keypoint location. However, when only a portion of the robot is visible, some keypoints may lie outside the image, leading to low confidence scores across the heatmap for these keypoints (Fig. 4). Selecting the pixel with the highest confidence in such cases can be misleading, as no pixel accurately represent the true keypoint. To address this, during evaluation, we apply a threshold $\epsilon$, only considering keypoints with confidence above it. For use with a PnP algorithm [20] for pose estimation, we require a minimum of four 2D-3D correspondences. If fewer than four keypoints remain after filtering, we iteratively reduce $\epsilon$ by 0.025 until at least four keypoints are retained.
Figure 4: The examples show predicted heatmaps for Joint 5 and the End-Effector overlaid on the original image. The End-Effector, being positioned outside the field of view, produces noisy heatmaps with lower confidence (measured by peak values). Heatmap pixel values are normalized for better visualization. The green arrows highlight the peak values for Joint 5 for visual clarity.
@@
where $P_{pred}^{i}$ and $P_{proj}^{i}$ are the $i^{th}$ keypoint in $P_{pred}$ and $P_{proj}$, respectively. During this sim-to-real finetuning, we use a lower learning rate than in the prior supervised training. To prevent model collapse, the Keypoint Net’s learning rate is set close to zero. Further details on learning rates are provided in the supplementary material.
@@
Region of Interest (RoI) Detection: During evaluation, we utilize the GroundingDINO [23] object detection model to automatically locate the region of interest around the robot. The detected region is cropped and resized to $224\times 224$ pixels. Further details are provided in the supplementary material. During training, we use ground truth bounding box information. However, to ensure our model’s robustness to region-of-interest detection, we employ a training curriculum: we use the ground truth bounding boxes and expand them by adding random offsets sampled from the uniform distribution $\mathcal{U}(0,\lambda)$ to their edges, with $\lambda$ progressing from 0 (first 30 epochs) to 30 pixels at epoch 30, 50 pixels at epoch 50, 80 pixels at epoch 70, 100 pixels at epoch 90, and 120 pixels at epoch 110. This approach encourages the model to generalize effectively, even with noisy region-of-interest detection during inference.
@@ 4.2.1 Robot Pose Prediction
RoboPEPP yields the highest scores across all sequences (except for Kuka DR where it remains competitive) among methods with unknown joint angles and bounding boxes. HPE [5], on the other hand, shows sensitivity to bounding box selection, with its performance dropping when ground truth bounding boxes are unavailable. Our analysis has shown that using bounding boxes just 5 pixels wider than ground truth reduces HPE’s accuracy by up to 25% on the Panda Photo test set and by around 50% with 10-pixel wider boxes. While both RoboPEPP and HPE are trained using ground truth bounding boxes, RoboPEPP’s training strategy (Sec. 3.3) reduces dependency on them. Notably, RoboPEPP outperforms HPE on most sequences even when HPE uses known bounding boxes during evaluation.
@@ 4.2.4 Percentage of Correct Keypoints
The accuracy of 2D keypoint detection affects the overall performance of RoboPEPP. Therefore, in Table 4, we report the percentage of correct keypoints (PCK) within thresholds of 2.5, 5, and 10 pixels on the Panda Photo dataset. Since HPE [5] and RoboPose [16] do not rely on 2D keypoint detection for pose estimation, we include only DREAM [19] as a baseline for comparison. RoboPEPP achieves high average PCK scores, with 0.43 @ 2.5 pixels, 0.73 @ 5 pixels, and 0.95 @ 10 pixels. While DREAM surpasses RoboPEPP in PCK@2.5 pixels on AK and XK, RoboPEPP achieves higher overall accuracy across most metrics, showcasing its robust keypoint detection. We hypothesize that DREAM’s known joint angles assumption allows it to focus entirely on keypoint detection enhancing PCK in some cases.
@@ 5 Conclusion
Sim-to-Real Self-Supervised Training: To bridge the sim-to-real gap, the trained networks are fine-tuned on real datasets with self-supervised training, as described in Sec. 3.3. An AdamW optimizer [24] is used with learning rates of $10^{-7}$ for the encoder and predictor and $10^{-5}$ for the Joint Network. The learning rate for the Keypoint Network is set close to zero to prevent model collapse. We observed that prioritizing Joint Network updates over the Keypoint Network yielded the best results. On the Panda RS dataset, for example, using a learning rate of $10^{-10}$ for the Keypoint Network and $10^{-5}$ for the Joint Network improved ADD AUC from 70.4 to 80.5. In contrast, reversing or equalizing the rates resulted in lower performance (71.9 and 76.4, respectively), highlighting the robustness of the keypoints for guiding the Joint Network. Further, gradients originating from the Keypoint Network’s output are scaled by $10^{-10}$ to prevent conflicting updates to the encoder-predictor. These learning rates are all decreased by a factor of $10^{8}$ over the training process. Models are fine-tuned separately for each real-world dataset for 10 epochs with a batch size of 64.
@@
We utilize the pre-trained GroundingDINO [23] object detection model to identify the region of interest, as described in Sec. 3.3. GroundingDINO is a highly accurate open-set object detector that accepts an (image, text) pair as input and outputs bounding boxes corresponding to regions of the image described by the text query. For all real and photo-realistic test datasets, we use the text query “robotic arm." However, for the Panda, Kuka, and Baxter domain-randomized datasets, we use the query “robot" because these images often contain multiple objects, some of which resemble arms and can confuse the detection model. All other parameters of GroundingDINO are left at their default values. To address scenarios where only a portion of the robot is detected, we expand all the detected bounding boxes, especially for real datasets. Increasing all the bounding box sizes by 100 pixels on all sides generally yields robust robot pose estimation results. However, some fine-tuning of this parameter may be necessary for optimal performance depending on the specific dataset. Nonetheless, high performance is obtained even without fine-tuning.
@@
We evaluate our method on the DREAM dataset [19], which includes sequences from three robots: Franka Emika Panda (Panda), Kuka iiwa7 (Kuka), and Rethink Robotics Baxter (Baxter). The dataset provides training and testing sequences in both synthetic and real-world settings, as detailed in Table A2. The synthetic data, created in Unreal Engine 4, comprises domain-randomized (DR) and photo-realistic (Photo) sequences. For real-world data, sequences of the Panda robot were captured using Microsoft Azure Kinect (AK), Xbox 360 Kinect (XK), and Intel RealSense D415 (RS) cameras, with the cameras positioned at fixed locations. Additionally, the Panda ORB dataset was collected using a RealSense camera but with varying camera placements. Example images from each dataset sequence are illustrated in Fig. A1.
@@
A5.3 Ablation: Joint Net and Keypoint Net
Fig. A3 and Table A3 analyze the Joint Net’s $G$ value (number of refinement steps) and Keypoint Net’s hidden channel dimension, respectively, with $G$=4 and 256 channels yielding the best overall results.
Figure A3: Average AUC and Joint Angle errors on the Panda dataset for different values of Joint Net’s G.
Channel Dim. AUC
128 69.3
256 75.5
512 74.4
Table A3: Average AUC on the Panda dataset for different values of the Keypoint Net’s hidden channel dimension.